Abstract Algebra. Please be detailed. Thank you.

+ 6x +3 E Z8[x]*. 3. Prove that 4x2
Answer #1

Solve het $(x) = 4x² + 6x + 3 € Z8 [x] then bix = 3(4x2 + 2x + 3) 2+ (2x) + 3) lo In Zorx 3(4x2) = 12x2 = 4x2 - 31(2x)3-17 28Please feel free to ask any query in the comment box and don't forget to rate if you like.